# Defects, parcellation, and renormalized negative diffusivities in non-homogeneous oscillatory media

### v1

@@ Abstract
The critical threshold $\gamma_{c}$ is finally derived by requesting $13\gamma_{c}\chi(0)^{2}=$ 1, which yields
$\gamma_{c}=\left[8\left(\sqrt{3}-\log\left(\frac{1+\sqrt{3}}{1-\sqrt{3}}\right)\right)\right]^{2}\simeq 11.027\,,$ (12)
@@
in excellent agreement with numerics, as shown in Fig. 4. Once the threshold is passed, $\varphi^{\prime}$ at the origin grows rapidly, which will squash the amplitude following $a=\sqrt{1-\gamma\varphi^{\prime 2}}$. The final stages of the process involve time-dependent effects ; it is still interesting to note that the renormalized diffusivity in Eq. (10) diverges when $a=\sqrt{1-\gamma\varphi^{\prime 2}}=0$, i.e., at the formation of the defect.
Our last step is to challenge the GL model to reproduce peristalsis data in Fig. 1. While “waxing and waning” and parcellation are easily captured, having the plateaus above the gradient at the corresponding spatial location is not captured by Eq. (1). However, as we noted, a gradient breaks parity and an additional term $V\nabla A$ is possible in Eq. (1), which represents advection to the right/left depending on the sign of $V$. Its effects will be detailed elsewhere ; the bottomline is that our results still apply but plateaus are advectively displaced and may thus be raised/lowered with respect to their Eulerian natural frequency counterparts. This allows us to reproduce all the features of experimental data as in Fig. 1.
@@
In conclusion, we have characterized the dynamics that leads to parcellation in non-homogeneous oscillatory systems. While our analysis specifically applies to a gradient of natural frequencies, as occurs in peristalsis, we expect that similar considerations apply to forced cases; see Ref. Mehta (2021) for forced GL. This is likely the case of vasomotion, where neural stimulation acts as a drive for the oscillations of arteriole vessels Haddock (2005); Drew (2020). Quantitative aspects of the neurovascular coupling remain mysterious, and will benefit from additional experiments. It would, in particular, be ideal to transiently shut down the neuronal activity and assess the undriven vasomotor dynamics. That will provide a sense of the elements that lead to the locking effects recently reported in Ref. Broggini (2024) and inform their modeling along the lines we presented here.
@@ I End Matter
Eq. (1) was integrated by a finite-difference scheme with fourth-order discretization of the Laplacian. We simulated half of the spatial interval and imposed symmetry conditions ($\mathrm{Re}$/$\mathrm{Im}[A]$ even/odd) at $x=0$. $N=601$ collocation points were used and the ratio of the time step $\Delta t$ to the smallest natural period satisfied $\frac{gL}{2\pi}\Delta t<10^{-2}$. Data analysis was performed in the stationary regime eventually reached in simulations.
@@
We identified plateaus by the clustering algorithm OPTICS Optics (2011). Its input is the matrix $R_{ij}=\frac{1}{T}\sum_{t=0}^{T}\left|e^{i(\phi_{i}(t)-\phi_{j}(t))}\right|\,$ that measures the $T$-averaged coherence of the phases $\phi_{i,j}(t)$ at positions $i,j$. Fig. 2 in NoteSM shows a typical instance of $R$. Three reasons motivated our OPTICS choice. First, it uses soft clustering, which leaves outliers (often corresponding to defects in our case) unclassified. Second, it is density-based, which is handy when the degree of synchronization is not uniform. Indeed, plateaus are not always “clean”, i.e., with sharp edges and vanishing frequency differences across them (see Akopov (2003); NoteSM ). Namely, extended plateaus alternating with smaller ones and/or unlocked regions may be observed (see Fig. 1 in NoteSM ). Third, the algorithm provides a systematic method to determine the number of clusters, which was important to identify the scaling (6).
WKB expansion.
We consider Eq. (1) for Dirichlet boundary conditions and locking at $\mp gL/2$ of two parcels symmetric around $x=0$, as in Fig. 2c. We pose $A\equiv\tilde{A}e^{-igL/2t}$ in the left region $-L/2\leq x\leq 0$, and focus on the bulk of the plateau where $|\tilde{A}|\ll 1$, i.e., away from $-L/2$ and the origin. Movie 1 in Ref. NoteSM suggests that the bulk profile of $\tilde{A}$ tends to become static and the unlocked region that hosts the defect at $x=0$ shrinks as $g$ increases. Eq. (1) in the bulk reduces then to
$0\simeq\frac{ig\left(x+\frac{L}{2}\right)}{D}\tilde{A}+\Delta\tilde{A}\,,$ (13)
where we neglected for simplicity the $r$ term with respect to the gradient. The general case is detailed in NoteSM .
The WKB method poses $\tilde{A}=e^{\psi}$, and expands $\psi=\psi_{0}+\psi_{1}+\ldots$ with the condition $\psi^{\prime\prime}\ll{\psi^{\prime}}^{2}$ Bender (1999). The resulting equations are
$(\psi_{0}^{\prime})^{2}\simeq-\frac{ig\left(x+\frac{L}{2}\right)}{D}\,,\qquad\psi_{1}^{\prime}\simeq\left(-\frac{1}{2}\log\psi^{\prime}_{0}\right)^{\prime}\,,$ (14)
whence Eq. (5) follows. The condition $\psi^{\prime\prime}\ll{\psi^{\prime}}^{2}$ requires $\left(x+\frac{L}{2}\right)\gg\ell_{p}=\left(D/|g|\right)^{1/3}$. The left-hand side being $\sim L$, it follows $\alpha\gg 1$. Neglecting $r$ with respect to $gx$ in Eq. (13) entails $r\ll|g|\ell_{p}$, i.e., $\beta\gg 1$ NoteSM .
@@
with $\nabla\rho(\zeta=\pm 1/2)=0$. Eq. (15) shows that its eigenvalues $\lambda_{n}$ are related to the $\sigma_{n}$’s of the Bloch-Torrey equation ($r^{\prime}=0$) by the shift $\lambda_{n}=\sigma_{n}+r^{\prime}$. We computed the $\sigma_{n}$’s numerically by using the ZGEEV function in LAPACK CIT (1999), which returns eigenvalues ordered by their real part. To span the $\alpha$ range, we fixed $D=0.5$, $L=1$, and varied $g=10^{0.1n}$, with the integer $n\in[0,60]$. Our numerics is consistent with Ref. Dyson (1991) with the change that $L=2$ instead of our $L=1$. To perform the verification, we re-derived results in Dyson (1991) as detailed in NoteSM .
@@
We consider the linearized GL Eq. (7) in the regime where the first branching occurred (see Fig. 3a,b). We show that branching is associated to strong modulation of the amplitude (precursor of defects) around the origin $x=0$. The pair of branched eigenvalues with the largest real part is denoted $\lambda_{1}^{R}\pm i\lambda_{1}^{I}$. The dominant asymptotics of the field $A$ at large times is
@@
where $\rho_{\pm}$ are the eigenmodes. Eq. (15) for the eigenmodes is symmetric under $\lambda^{I}\mapsto-\lambda^{I}$, $x\mapsto-x$, $\rho\mapsto\rho^{*}$, which implies $\rho_{-}\left(x\right)=\rho_{+}^{*}\left(-x\right)$ Dyson (1991). The expression (16) at the origin reduces then to $2e^{\lambda_{1}^{R}t}|\rho_{+}\left(0\right)|\cos\left(\lambda_{1}^{I}t+arg\left(\rho_{+}\left(0\right)\right)\right)$, which vanishes periodically with period $2\pi/|\lambda_{1}^{I}|$.
@@ Acknowledgements.
We thank Jacob Duckworth, Vladimir Matchkov, and Marcelo Rozenberg for relevant discussions. This work was supported by the NIH U19 Cooperative Agreement Grant (U19 NS123717).

### v2

@@ Abstract
The critical threshold $\gamma_{c}$ is finally derived by imposing $3\gamma_{c}\chi(0)^{2}=1$, which yields
$\gamma_{c}=\left[8\left(\sqrt{3}-\log\left(\frac{\sqrt{3}+1}{\sqrt{3}-1}\right)\right)\right]^{2}\simeq 11.027\,$, (11)
@@
in agreement with numerics (see Figs. 3a-b). Once the threshold is passed, $\varphi^{\prime}$ at the origin grows rapidly, which reduces the amplitude as $a=\sqrt{1-\gamma\varphi^{\prime 2}}$. The final stages of the process involve time-dependent effects ; it is still interesting to note that the renormalized diffusivity in Eq. (9) diverges when $a=\sqrt{1-\gamma\varphi^{\prime 2}}=0$, i.e., at the formation of the defect.
Knowledge of the critical curve allows us to determine the type of transitions taking place along it. Equations (7) and (8) for the locked case have their left-hand sides vanishing (see NoteSM ). The amplitude is indeed stationary and even in space, and the phases increase with a common frequency $\omega$, which vanishes as $\varphi(x)$ is odd. The no-flux boundary conditions leave unknown the boundary amplitudes that satisfy the constraint a(L/2)=a(L/2)a(-L/2)=a(L/2). Solutions, i.e., the values of a(L/2)a(-L/2) reported in Fig. 3c-d, are sought by a shooting method and the software XPPAUT xppaut (2003), which allows to follow solutions as parameters vary. Figure 3 summarizes the conclusions : two locked solutions are present, one stable (selected by the dynamics) and one unstable, that merge at the boundary between nonlinear phase locking and plateaus. The presence of two solutions and their merging indicates the saddle-node nature of the bifurcations along the critical curve.
Figure 4: (Color online) Properties of plateaus. The number of plateaus $N_{p}$ is shown vs $\alpha^{1/3}$ for Dirichlet (circles) and no-flux (squares) boundary conditions. Empty/full symbols refer to $\beta=10^{-2}$ and $\beta=10^{-4}$, respectively. The insets show plateaus for no-flux boundary conditions with $L=1$, $\beta=10^{-4}$ and (a) $\alpha=1002$ or (b) $\alpha=252$.
In the last region of the phase diagram (orange in Fig. 2), at least a pair of eigenmodes grows linearly and stabilizes by nonlinear effects into multiple plateaus (see Fig. 4). The simplest geometry has two sharp plateaus, as in Fig. 4c. Each plateau features an average frequency, i.e., $\omega_{1}$ and $\omega_{2}$, respectively, common to all its oscillators. That ensures that their phases vary in time with the same dominant term yet may not keep a fixed phase difference as in strict phase locking Pikovsky (2001). The jump $\Delta\omega=|\omega_{1}-\omega_{2}|$ dictates a characteristic time $2\pi/\Delta\omega$ for the two plateaus to accumulate phase differences. As phase gradients increase at the boundary of the plateaus, the amplitude damping term in Eq. (2) drives down the amplitude. When the difference of the phases in the two plateaus reaches $2\pi$, a defect $A=0$ is produced, the phase is undefined, and the $2\pi$ difference is reset to restart the periodic process (see Movie 1 in Ref. NoteSM ). As already noted, increasing $\alpha$ reduces the range of diffusion and increases the number of plateaus $N_{p}$, as shown in Fig. 4. Numerical results identify a clean scaling $N_{p}\propto\frac{L}{\ell_{p}}=\alpha^{1/3}$, which is independent of boundary conditions as it is also observed for Dirichlet boundary conditions $A\left(\mp L/2\right)=e^{\mp igL/2}$. To understand the $1/3$ scaling, we observe that uncoupled oscillators spaced by $\ell$ accumulate phase differences at a rate $|g|\ell$, reaching differences $\simeq 2\pi$ over times $\tau_{\ell}\propto(|g|\ell)^{-1}$. Diffusive couplings extend their spatial range as the square-root of time. For two points to be coupled strongly enough to lock, it is necessary that $\ell\propto\sqrt{D\tau_{\ell}}$. We conclude that the typical length of the plateaus is $\ell_{p}\equiv\left(\frac{D}{|g|}\right)^{1/3}$ and
$N_{p}\propto\frac{L}{\ell_{p}}=\alpha^{1/3}\,.$ (12)
The scaling (12) is consistent with each point oscillating independently ($\ell_{p}\to 0$) in the uncoupled $D\to 0$ limit. The range of validity of the scaling (12) depends on $\beta$, defined in Eq. (5). As $\beta$ increases for a fixed large $\alpha$ (or $\alpha$ decreases), $N_{p}$ is reduced, as confirmed in a limited range in Fig. 4, and globally by the $(\alpha,\beta)$ supplementary plots of Fig. S5 NoteSM , until the scaling (12) breaks down.
Our last step is to challenge the GL model to reproduce peristalsis data in Fig. 1. While “waxing and waning” and parcellation are easily captured, having the plateaus above the frequencies of the uncoupled system at the corresponding spatial locations is not captured by Eq. (1). However, as we noted, Eq. (1) with real coefficients is not the most general form consistent with the presence of a gradient. First, since parity is broken by $g$, an additional term $V\nabla A$ is possible by symmetry. Such term represents advection to the right/left depending on the sign of $V$, which may occur naturally in the physiological examples mentioned in the introduction. Indeed, peristalsis Diamant (1969); Parsons (2020) is associated with aboral flows, and vasomotion with the cortical flow of blood Haddock (2005); Drew (2020); Broggini (2024). Advection displaces the plateaus, which may thus be raised/lowered with respect to their Eulerian counterparts. Second, both $r$ and $D$ may be complex. Complex $r$’s are known to shift the frequency of oscillation of the Stuart-Landau equation Pikovsky (2001) but they are not associated to the coupling among oscillators. In other words, they cannot account for the differences observed in Fig. 1 between the intact and the post hoc sectioning of the intestine Diamant (1969). Conversely, complex $D$’s do involve coupling among different oscillators and can raise/lower the frequency of the plateaus. Allowing for a combination of complex $D$ and advection leads to the graphs in Fig. 1, which capture the main experimental features. Since the same holds for graphs obtained individually with advection or complex $D$ NoteSM , their disentanglement calls for additional experiments.
@@
In conclusion, we characterized the dynamics that leads to parcellation in non-homogeneous oscillatory systems. While our analysis specifically applies to a gradient of natural frequencies, as in peristalsis, we expect that similar considerations apply to forced cases Diamant (1969); Parsons (2020); see Ref. Mehta (2021) for forced GL. This is likely the case of vasomotion, where neural stimulation acts as a drive for the oscillations of arteriole vessels Drew (2020); Broggini (2024). Quantitative aspects of the neurovascular coupling remain mysterious, and will benefit from additional experiments. It would, in particular, be ideal to transiently shut down the neuronal activity and assess the undriven vasomotor dynamics. That will provide a sense of the elements that lead to locking effects, as recently reported in Ref. Broggini (2024), and inform their modeling along the lines presented here.
@@ I End Matter
We integrated Eq. (1) by a finite-difference scheme with fourth-order discretization of the Laplacian. We simulated half of the spatial interval and imposed symmetry conditions ($\mathrm{Re}$$\mathrm{Im}[A]$ even/odd) at $x=0$. We used $N=601$ collocation points, and the ratio of the time step $\Delta t$ to the smallest natural period satisfied $\frac{gL}{2\pi}\Delta t<10^{-2}$. Data analysis was performed in the stationary regime eventually reached in simulations.
@@
We identified plateaus by the clustering algorithm OPTICS Optics (2011). Its input is the matrix $R_{ij}=\frac{1}{T}\sum_{t=0}^{T}\left|e^{i(\phi_{i}(t)-\phi_{j}(t))}\right|\,$ that measures the $T$-averaged coherence of the phases $\phi_{i,j}(t)$ at positions $i,j$ (see Fig. S3 in NoteSM ). Three reasons motivated our choice of OPTICS. First, classification is not forced in uncertain cases and outliers, which often correspond to defects in our case, are left unclassified. Second, it is density-based, which is handy for non-uniform extents of synchronization. Indeed, plateaus are not always “clean”, i.e., with sharp edges and vanishing frequency differences along them Akopov (2003); NoteSM , so that extended plateaus alternating with smaller ones and/or unlocked regions may be observed (see Fig. S2 in NoteSM ). Third, the algorithm provides a systematic method to determine the number of clusters, which was important to identify scaling (12).
@@
with $\nabla\rho(\zeta=\pm 1/2)=0$. Eq. (13) shows that its eigenvalues $\lambda_{n}$ are related to the $\sigma_{n}$’s of the Bloch-Torrey equation ($r^{\prime}=0$) by the shift $\lambda_{n}=\sigma_{n}+r^{\prime}$. We computed the $\sigma_{n}$’s by using a second-order discretization of the Laplacian in Eq. (4) and using the ZGEEV function in LAPACK CIT (1999) to compute the eigenvalues. To span the $\alpha$ range, we fixed $D=0.5$, $L=1$, and varied $g=10^{0.1n}$, with the integer $n\in[0,60]$. Our numerics is consistent with Ref. Dyson (1991) with the change that $L=2$ instead of our $L=1$. Results of Dyson (1991) are re-derived in SM NoteSM .
@@
We consider the linearized GL Eq. (4) in the regime where the first branching occurred (see Fig. 3a,b). We show that branching is associated with strong modulation of the amplitude (precursor of defects) around $x=0$. The pair of branched eigenvalues with the largest real part is denoted $\lambda_{1}^{R}\pm i\lambda_{1}^{I}$. The dominant asymptotics of the field $A$ at large times is
@@
where $\rho_{\pm}$ are the eigenmodes. Eq. (13) for the eigenmodes is symmetric under $\lambda^{I}\mapsto-\lambda^{I}$, $x\mapsto-x$, $\rho\mapsto\rho^{*}$, which implies $\rho_{-}\left(x\right)=\rho_{+}^{*}\left(-x\right)$ Dyson (1991). The expression (14) at the origin reduces to $2e^{\lambda_{1}^{R}t}|\rho_{+}\left(0\right)|\cos\left[\lambda_{1}^{I}t+{\rm arg}\left(\rho_{+}\left(0\right)\right)\right]$, which vanishes periodically with period $2\pi/|\lambda_{1}^{I}|$.
@@ Acknowledgements.
We thank Jacob Duckworth, Vladimir Matchkov, and Marcelo Rozenberg for relevant discussions. This work was supported by the NIH grants U19 NS123717 and U19 NS137920.